Carlos earned a total of $225,000 for 2018. How much in FICA tax will his employer be required to withhold in his name?
Answer #1
Social Security tax 7960.80 =128400*6.2%
Medicare tax 3262.50 =225000*1.45%
Additional Medicare tax 225.00 =(225000-200000)*0.9%
FICA tax to be withheld 11448.30
The 2018 wage limit for Social Security taxes is $128400
Additional Medicare tax is payable for wages above $200000